OpenStudy (anonymous):

Steam at 100bar,375 degrees celcitus expands isentropically in a cylinder behind a piston to a pressure of 10 bar.Calculateper kilogram of steam the workdone

OpenStudy (anonymous):

As I understands it you then get a loss of 90 bar Pascal is more used as pressure in thermodynamics math. The bar is defined using the SI unit pascal, namely: 1 bar ≡ 100,000 Pa. 1 bar is therefore equal to: 100 kPa or 1×105 N/m2 And the thermodynamics: http://hyperphysics.phy-astr.gsu.edu/hbase/thermo/firlaw.html hope this helps

OpenStudy (anonymous):

The absolute temperature in a gas is given by the average kinetical energy \[E _{k}\] of them molecules in the gas given by the formula \[E _{k}=\frac{ 3 }{ 2 }k \times T\] \[k = 1,38 \times 10^{-23} J \times K \] Boltzmanns constant. Now you should have a better chance of getting this assignment done.